$SPAD/src/input segbind.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool segbind.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
x = a..b
--R
--R
--R   (1)  x= a..b
--R                                                 Type: SegmentBinding(Symbol)
--E 1
--S 2 of 6
sum(i**2, i = 0..n)
--R
--R
--R          3     2
--R        2n  + 3n  + n
--R   (2)  -------------
--R              6
--R                                          Type: Fraction(Polynomial(Integer))
--E 2
--S 3 of 6
draw(x**2, x = -2..2)
--R
--I   Compiling function %B with type DoubleFloat -> DoubleFloat
--R   Graph data being transmitted to the viewport manager...
--R   AXIOM2D data being transmitted to the viewport manager...
--R
--R   (3)  TwoDimensionalViewport: "x^2"
--R                                                 Type: TwoDimensionalViewport
--E 3
--S 4 of 6
sb := y = 1/2..3/2
--R
--R
--R            1    3

--R   (4)  y= (-)..(-)
--R            2    2
--R                                      Type: SegmentBinding(Fraction(Integer))
--E 4
--S 5 of 6
variable(sb)
--R
--R
--R   (5)  y
--R                                                                 Type: Symbol
--E 5
--S 6 of 6
segment(sb)
--R
--R
--R         1    3
--R   (6)  (-)..(-)
--R         2    2
--R                                             Type: Segment(Fraction(Integer))
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5